if the K.E of electron is Increased by 4 times then Broglie wavelength associated with it would become????????


 

Grade:11

2 Answers

vikas askiitian expert
13 years ago

de broglie wavelength (lamda)L=  h/(2mKE)1/2                 ..............1          

 

 here h is planks constant, m is mass of electron  & KE is kinetic energy...

 

     from eq 1

                      L(lamda) is inversly proportional to (KE)1/2

                    L/L1   =   {(KE)f /(KE)i}1/2          

(KE)f = 4(KE)i   (given)

 so    L/L1 = 2   or

        L1 = L/2

 it means de broglie wavelength of electron is decreased to half times the  initial value....

 

PRASHANT MISHRA
13 years ago

According to De BROGLIE,

wavelength=h/root 2Em

where h....Planck constant

E.... Kinetic energy, m= mass of electron

Thus, when kinetic energy associated with electron is increased four times, then wave length would become half of the initial value.